Brest State Technical University

(http://www.bstu.unibel.by/index-e.html)

Information

Brest State Technical University is the largest scientific and educational center in the western part of republic Belarus. Training of highly qualified specialists as well as fundamental scientific research work is carried out in the field of construction, architecture, electronics, mechanical engineering, economy and ecology.

General Information

The scientific potential of the University includes 12 Doctors of Science, 145 Candidates (Ph.D.) and experienced academic teachers. The training of 4000 full-time students and 1400 part-time students is conducted at 5 departments:

• Civil Engineering;

• Mechanical Engineering and Electronics;

• Water Supply Systems & Soil Conservation;

• Economical;

• Extra-Mural Studies.

Auxiliary functions are fulfilled by: Preparatory Department and The Institute of further Education and Retraining.

The training course lasts 4 years and 10 months for full-time students while 5 years and 10 months for part-time students.

Brief History

This University began as a Civil Engineering Institute on April 1, 1966. The first intake was 330 full-time students and 110 evening-class students. The teaching staff numbered 32 teachers. In 1969 the number of students reached 2700, namely 1960 full-time students, 480 evening-class students, 260 part-time students. The teaching staff increased till 186 teachers. In 1989 the institute was reorganized into Polytechnic Institute. Since then Mechanical Engineering and Electronics Department and Economical Department were opened, new specialties appeared; the spectrum of research work has expanded. Now it is the largest technical institution of higher learning in the western region of Belarus. Since its foundation more than 18000 specialists have graduated from this Institution, which was incorporated as a State Technical University in 2000.

Development Tendencies

The development of the University is adapted to satisfy the needs of the Belarus Republic and of Brest region in specialists: The conditions are being created for highly-qualified training of economists and managers; The range of new specialties connected with electronics and computers is being expanded; The Scientific Research Institute for the problems of Construction Engineering organized in May 2004 is successfully being developed; The process of reformation of the system of the University is being carried out to offer Master and Bachelor programs.

International Contacts

Our University develops close contacts with higher educational establishments in Germany, Poland, Ukraine and Russia. We have long-term partner contacts with Bialystock Polytechnic Institute (Poland), Lublin Polytechnic Institute (Poland), Higher Technical Professional School in Biberach(Germany) and Higher Technical School in Ravensburg-Weingarten. This partnership creates an essential basis for mutually useful training activity and scientific researches.

Contact Information

Brest State Technical University 267, Moskovskaja Str., Brest, Belarus, 224017 Phone: + 375 162 42-33-93 + 375 162 42-40-84 Fax: + 375 162 42-21-27 E-mail: canc@bstu.by